# Machine Learning for Synthetic Data Generation: A Review

### v9

@@ I Introduction
Nonetheless, it is important to recognize that machine learning does not invariably resolve problems or yield the optimal solution. Despite artificial intelligence is currently experiencing a golden age, numerous challenges persist in the development and application of machine learning technology [4]. As the field continues to advance, addressing these obstacles will be essential for unlocking the full potential of machine learning and its transformative impact on various industries. Machine learning endows intelligent computer systems with the capacity to autonomously tackle tasks, pushing the envelope of industrial innovation [1]. By integrating high-performance computing, contemporary modeling, and simulations, machine learning has evolved into an indispensable instrument for managing and analyzing massive volumes of data [2, 3].
@@
Addressing these challenges will be essential for unlocking the full potential of machine learning and its transformative impact on various industries [9, 10, 11]. Generally, synthetic data are defined as the artificially annotated information generated by computer algorithms or simulations [12, 4]. Synthetic data is generally defined as artificially annotated information generated by computer algorithms or simulations [12, 4]. In many cases, synthetic data is necessary when real data is either unavailable or must be kept private due to privacy or compliance risks [10, 13, 14]. This technology is extensively utilized in various sectors, such as healthcare, business, manufacturing, and agriculture, with demand growing at an exponential rate [15].
@@
  • We examine a diverse array of deep neural network architectures and deep generative models dedicated to generating high-quality synthetic data, present advanced generative models, and outline potential avenues for future research (Section III and IV).
  • We address privacy and fairness concerns, as sensitive information can be inferred from synthesized data, and biases embedded in real-world data can be inherited. We review current technological advancements and their limitations in safeguarding data privacy and ensuring the fairness of synthesized data (Section V and VI).
  • We outline several general evaluation strategies to assess the quality of synthetic data (Section VII).
  • We identify challenges faced in generating synthetic data and during the deployment process, highlighting potential future work that could further enhance functionality (Section VIII).
@@ II-C Natural Language Processing (NLP)
Despite these advancements, challenges remain. For example, the meaning of words and phrases can change depending on their context, and homonyms with distinct definitions can pose additional difficulties [65]. To tackle these challenges, the BLEURT model was proposed, which models human judgments using a limited number of potentially biased training examples based on BERT. The researchers employed millions of synthetic examples to develop an innovative pre-training scheme, bolstering the model’s ability to generalize [66]. Experimental results indicate that BLEURT surpasses its counterparts on both the WebNLG Competition dataset and the WMT Metrics, highlighting its efficacy in NLP tasks [39].
@@ II-D Healthcare
In order to protect health information and improve reproducibility in research, synthetic data has drawn mainstream attention in the healthcare industry [68, 69]. Many labs and companies have harnessed the tools of big data and advanced computation tools to produce large quantities of synthetic data [70]. Modeled after patient data, synthetic data generation is essential to understanding diseases while maintaining patient confidentiality and privacy simultaneously [71]. Theoretically, synthetic data can reflect the original distribution of the data instead of revealing actual patient data [71, 72, 73].
@@
In healthcare, patient information is often stored in electronic health records (EHR) format.  [86, 87, 88]. Research in medicine has been greatly facilitated by the availability of information from electronic health records [89, 90]. MedGAN, an adversarial network model for generating realistic synthetic patient records, has been proposed by Edward Choi and other colleagues. With the help of an autoencoder and generative adversarial networks, medGAN can generate high-dimensional discrete variables (e.g., binary and count features) based on real patient records [16]. Based on their evaluations of medGAN’s performance on a set of diverse tasks reported, including reporting distribution statistics, classification performance [91], and expert review, medGAN exhibits close-to-real-time performance [92, 93, 94, 95, 16]. Using synthetic data can help reduce the regulatory barriers preventing the widespread sharing and integration of patient data across multiple organizations in the past [96, 97]. Researchers across the globe would be able to request access to synthetic data from an institution to conduct their own research using the data. Such capabilities can increase both the efficiency and scope of the study as well as reduce the likelihood of biases being introduced into the results [69, 98, 99].
@@ II-I Finance
Synthetic data generation also transforms the financial services industry by enabling more accurate risk assessments and fraud detection[102]. Synthetic data generation can identify anomalies and potential risks by simulating financial transactions and market behaviors, allowing financial institutions to implement more effective fraud prevention measures and develop more resilient financial strategies. Furthermore, synthetic data generation can support compliance with regulatory requirements by providing detailed, real-time reporting and analysis of financial activities [47]. In the context of human resources, synthetic data generations can model workforce dynamics, including employee performance, engagement, and turnover. By analyzing these models, businesses can develop strategies to improve employee satisfaction, enhance productivity, and reduce turnover rates. For example, synthetic data generations can simulate the impact of various HR policies on workforce morale and performance, helping HR departments to implement the most effective practices.
@@ II-J Other Applications
III Deep Neural Network
It is no secret that deep neural networks have become increasingly prominent in the field of computer vision and in other areas. Nevertheless, they require large amounts of annotated data for supervised training, which limits their effectiveness [137]. In this section, we review and compare various commonly-used deep neural network architectures as background knowledge, including multiple layer perception (MLP) in Section III-A, convolutional neural network (CNN) in Section III-B, recurrent neural network (RNN) in Section III-C, graph neural network (GNN) in Section III-D and transformer in Section III-E.
III-A Multiple Layer Perception (MLP)
Multiple-layer perception (MLP) is a classic feedforward artificial neural network that uses a fully connected connection in a single layer. It models all the possible interactions between the features. It is also the most used neural network. In the $i$-th layer, the propagation can be written as
$h^{i+1}=\sigma(\mathbf{W}^{T}h^{i}+b),$ (1)
where $h^{i}$ is the input and $h^{i+1}$ is the output (which is also the input of the $i+1$-th layer), $\mathbf{W}$ is weight matrix, $b$ is the bias vector, both $\mathbf{W}$ and $b$ are parameters of MLP. $\sigma()$ denotes the activation function. Popular activation functions incorporate the sigmoid, ReLU, tanh, etc. The goal of the activation function is to provide nonlinear transformation. MLP is the basis of many neural network architectures. Please refer to [138] for more details about MLP.
III-B Convolutional Neural Network (CNN)
Convolutional neural network (CNN) was proposed to learn a better representation of images [138]. The core idea of the convolutional neural network is to design a two-dimensional convolutional layer that slides over the image to model the small-sized patch horizontally and vertically. Convolutional neural networks with one-dimensional convolutional layers can also model sequence data. Please refer to [138] for more details about CNN.
III-C Recurrent Neural Network (RNN)
The sequence is one of the most popular data formats in real-world applications, such as natural language and speech signals. Recurrent neural network (RNN) was designed to represent and generate sequence data [139]. Suppose we have a sequence of length $T$, $X=[x_{1},x_{2},...,x_{T}]$, $x_{t}$ denotes the $t$-th token, recurrent neural network designs a recursive (or recurrent) mechanism to reuse the neural network module. Specifically, at the $t$-th step, we have
$o_{t},h_{t}=\text{RNN}(x_{t},h_{t-1}),$ (2)
where $h_{t-1}$ denotes the hidden state at the $t-1$-th step and is used to represent all the historical memory from previous $t-1$ tokens ($x_{1},x_{2},...,x_{t-1}$); $o_{t}$ denotes the output of the RNN at the $t$-th step. The RNN module is recursively reused for each token. However, the vanilla RNN structure suffers from gradient vanishment issues, especially for long sequences. The state-of-the-art RNN architectures long short-term memory (LSTM) [140] and gated recurrent unit (GRU) [141] design a specialized gate to save the memory to alleviate the issue.
III-D Graph Neural Network (GNN)
Many graph-structured data exist in downstream applications, such as social networks, brain networks, chemical molecules, and knowledge graphs. Graph neural network was proposed to model graph-structured data and learn the topological structure from graph [142, 143, 144]. Specifically, graph neural networks build the interaction between the connected nodes and edges to model the topological structure of the graph [145, 146]. The feedforward rule of graph neural network can be formulated as
$h_{i}^{(l+1)}=\text{Aggregate}_{j\in\mathcal{N}(i)}(h_{j}^{(l)},m_{j},m_{ji}),$ (3)
$l=1,\cdots,L,$
where $h_{i}^{(l)}$ denotes the embedding of the $i$-th node at the $l$-th layer, $m_{j}$ denotes the node feature of the node $j$, $m_{ji}$ denotes the edge feature of the edge that connects nodes $j$ and $i$, $\mathcal{N}(i)$ is the set of nodes that connect with $i$. Within each layer, we update the current representation via aggregating the information from (1) representation from the previous layer, (2) node feature (3) edge feature. For example, in chemical compounds, each node corresponds to an atom, and the node feature is the category of the atom, e.g., Carbon atom, Nitrogen atom, Oxygen atom; each edge is a chemical bond, and the edge feature is the type of the bond, including single bond, double bond, triple bond, and aromatic bond.
III-E Transformer
The Transformer architecture, introduced by Vaswani et al. in the groundbreaking paper “Attention is All You Need” in 2017, revolutionized the field of natural language processing and machine learning. Unlike traditional sequential models like RNNs (Recurrent Neural Networks) and LSTMs (Long Short-Term Memory), the Transformer model leverages a unique mechanism called self-attention, which allows it to capture long-range dependencies and relationships within the input data more effectively. This architecture consists of a stack of identical layers, each containing a multi-head self-attention mechanism followed by position-wise fully connected feed-forward networks. By eschewing recurrent or convolutional layers, the Transformer model is highly parallelizable and computationally efficient, leading to faster training times and improved performance on various NLP tasks.
The Generative Pre-trained Transformer (GPT) is a cutting-edge deep learning model that has revolutionized natural language processing (NLP) tasks [147]. Developed by OpenAI, GPT is an autoregressive transformer-based model that has displayed unparalleled performance in tasks such as text generation, translation, text summarization, and question-answering.
The model’s architecture consists of multiple self-attention mechanisms and position-wise feedforward layers, enabling it to capture long-range dependencies and generate highly coherent and contextually relevant text. The key to GPT’s success lies in its unsupervised pre-training on vast amounts of textual data, followed by fine-tuning on specific tasks. As the GPT model series progresses, with GPT-3 being the latest version at the time of this writing, the size and capabilities of the model continue to grow, paving the way for increasingly sophisticated NLP applications and opening up new possibilities for the generation of synthetic data.
By leveraging its pre-training on massive datasets and fine-tuning for specific tasks, GPT can produce artificial data samples that closely resemble real-world data. This capability is particularly valuable in scenarios where access to real data is limited due to privacy, regulatory, or resource constraints. GPT-generated synthetic data can be used to augment existing datasets, enabling researchers and practitioners to build more robust and accurate machine learning models while mitigating the risks associated with using sensitive or private data. Additionally, the synthetic data generated by GPT models can help address challenges related to data scarcity, class imbalance, or the need for domain-specific data, ultimately contributing to developing and deploying more effective AI solutions across various domains and applications.
IV Generative AI
@@
Generative AI models refer to a wide class of AI methods that could learn the data distribution from existing data objects and generate novel structured data objects, which fall into the category of unsupervised learning. Generative AI models, also known as deep generative models, or distribution learning methods, learn the data distribution and samples from the learned distribution to produce novel data objects. In this section, we investigate several generative AI models that are frequently used in synthetic data generation, including the language model in Section IV-A, variational autoencoder (VAE) in Section IV-C, generative adversarial network (GAN) in Section IV-D, reinforcement learning (RL) in Section IV-E, and diffusion model in Section IV-F. Table II compares various generative AI methods from several aspects.
@@ IV-A Language Model
IV-B Self-Supervised Learning (SSL)
@@ IV-C Variational Autoencoder (VAE)
IV-D Generative Adversarial Network (GAN)
@@ IV-F Diffusion Model
IV-G Multimodal Learning
Multimodal data refers to datasets that integrate multiple types of data, such as text, images, audio, and numerical values. This type of data provides a comprehensive view by combining different sources of information, which is crucial for tasks requiring a holistic understanding of complex scenarios. In fields like healthcare [167], finance, and autonomous systems, multimodal data enables more accurate and robust analysis and decision-making by leveraging the strengths of each data type. For instance, in drug discovery, multimodal data can combine genomic data, chemical structures, and clinical outcomes to enhance the prediction of drug efficacy and safety [93].
@@ IV-G Multimodal Learning
V Privacy Risks and Prevention
@@
V-A Privacy Risks in Data Synthesis
@@
V-B Privacy Protection in Data Synthesis
@@
V-C Privacy Threats in Foundation Models
@@ VII Evaluation Strategy
    Evaluation using a pre-trained machine learning model. As mentioned in Section IV-D, in the generative adversarial network (GAN), the discriminator differentiates fake data (synthesized data) from real ones. Consequently, the output of the discriminator can measure how closely synthetic data resembles real data. The performance of the discriminator on the synthesized data can be used as an indicator of how well the generator produces realistic data. This strategy can be applied not only to GANs but also to other generative models where a pre-trained machine learning model is used for evaluation.

### v10

@@ I Introduction
Nonetheless, it is important to recognize that machine learning does not invariably resolve problems or yield the optimal solution. Despite artificial intelligence is currently experiencing a golden age, numerous challenges persist in the development and application of machine learning technology [4]. As the field continues to advance, addressing these obstacles will be essential for unlocking the full potential of machine learning and its transformative impact on various industries.
@@
Tackling these issues is crucial to fully realizing the transformative power of machine learning across diverse sectors [9, 10, 11]. Generally, synthetic data are defined as the artificially annotated information generated by computer algorithms or simulations [12, 4]. In many cases, synthetic data is necessary when real data is either unavailable or must be kept private due to privacy or compliance risks [10, 13, 14]. This technology is extensively utilized in various sectors, such as healthcare, business, manufacturing, and agriculture, with demand growing at an exponential rate [15].
@@
  • We examine a diverse array of deep generative models dedicated to generating high-quality synthetic data, present advanced generative models, and outline potential avenues for future research (Section III).
  • We address privacy and fairness concerns, as sensitive information can be inferred from synthesized data, and biases embedded in real-world data can be inherited. We review current technological advancements and their limitations in safeguarding data privacy and ensuring the fairness of synthesized data (Section IV and V).
  • We outline several general evaluation strategies to assess the quality of synthetic data (Section VI).
  • We identify challenges faced in generating synthetic data and during the deployment process, highlighting potential future work that could further enhance functionality (Section VII).
@@ II-C Natural Language Processing (NLP)
Despite these advancements, challenges remain. For example, the meaning of words and phrases can change depending on their context, and homonyms with distinct definitions can pose additional difficulties [65]. To tackle these challenges, the BLEURT model was proposed, which models human judgments using a limited number of potentially biased training examples based on BERT. The researchers employed millions of synthetic examples to develop an innovative pre-training scheme, bolstering the model’s ability to generalize [66, 67]. Experimental results indicate that BLEURT surpasses its counterparts on both the WebNLG Competition dataset and the WMT Metrics, highlighting its efficacy in NLP tasks [39].
@@ II-D Healthcare
In order to protect health information and improve reproducibility in research, synthetic data has drawn mainstream attention in the healthcare industry [69, 70]. Many labs and companies have harnessed the tools of big data and advanced computation tools to produce large quantities of synthetic data, or digital twin [71, 72]. Modeled after patient data, synthetic data generation is essential to understanding diseases while maintaining patient confidentiality and privacy simultaneously [73]. Theoretically, synthetic data can reflect the original distribution of the data instead of revealing actual patient data [73, 74, 75].
@@
In healthcare, patient information is often stored in electronic health records (EHR) format [88, 89, 90]. Research in medicine has been greatly facilitated by the availability of information from electronic health records [91, 92]. MedGAN, an adversarial network model for generating realistic synthetic patient records, has been proposed by Edward Choi and other colleagues. With the help of an autoencoder and generative adversarial networks, medGAN can generate high-dimensional discrete variables (e.g., binary and count features) based on real patient records [16]. Based on their evaluations of medGAN’s performance on a set of diverse tasks reported, including reporting distribution statistics, classification performance [93], and expert review, medGAN exhibits close-to-real-time performance [94, 95, 96, 97, 16]. Using synthetic data can help reduce the regulatory barriers preventing the widespread sharing and integration of patient data across multiple organizations in the past [98, 99]. Researchers across the globe would be able to request access to synthetic data from an institution to conduct their own research using the data. Such capabilities can increase both the efficiency and scope of the study as well as reduce the likelihood of biases being introduced into the results [70, 100, 101].
@@ II-I Finance
Synthetic data generation also transforms the financial services industry by enabling more accurate risk assessments and fraud detection[104]. Synthetic data generation can identify anomalies and potential risks by simulating financial transactions and market behaviors, allowing financial institutions to implement more effective fraud prevention measures and develop more resilient financial strategies. Furthermore, synthetic data generation can support compliance with regulatory requirements by providing detailed, real-time reporting and analysis of financial activities [47]. In the context of human resources, synthetic data generations can model workforce dynamics, including employee performance, engagement, and turnover. By analyzing these models, businesses can develop strategies to improve employee satisfaction, enhance productivity, and reduce turnover rates. For example, synthetic data generation can simulate the impact of various HR policies on workforce morale and performance, helping HR departments to implement the most effective practices.
@@ II-J Other Applications
III Generative AI
@@
Generative AI models refer to a wide class of AI methods that could learn the data distribution from existing data objects and generate novel structured data objects, which fall into the category of unsupervised learning. Generative AI models, also known as deep generative models, or distribution learning methods, learn the data distribution and samples from the learned distribution to produce novel data objects. In this section, we investigate several generative AI models that are frequently used in synthetic data generation, including the language model in Section III-A, variational autoencoder (VAE) in Section III-C, generative adversarial network (GAN) in Section III-D, reinforcement learning (RL) in Section III-E, and diffusion model in Section III-F. Table II compares various generative AI methods from several aspects.
@@ III-A Language Model
III-B Self-Supervised Learning (SSL)
@@ III-C Variational Autoencoder (VAE)
III-D Generative Adversarial Network (GAN)
@@ III-F Diffusion Model
III-G Multimodal Learning
Multimodal data refers to datasets that integrate multiple types of data, such as text, images, audio, and numerical values. This type of data provides a comprehensive view by combining different sources of information, which is crucial for tasks requiring a holistic understanding of complex scenarios. In fields like healthcare [158], finance, and autonomous systems, multimodal data enables more accurate and robust analysis and decision-making by leveraging the strengths of each data type. For instance, in drug discovery, multimodal data can combine genomic data, chemical structures, and clinical outcomes to enhance the prediction of drug efficacy and safety [95, 159].
@@ III-G Multimodal Learning
IV Privacy Risks and Prevention
@@
IV-A Privacy Risks in Data Synthesis
@@
IV-B Privacy Protection in Data Synthesis
@@
IV-C Privacy Threats in Foundation Models
@@ VI Evaluation Strategy
    Evaluation using a pre-trained machine learning model. As mentioned in Section III-D, in the generative adversarial network (GAN), the discriminator differentiates fake data (synthesized data) from real ones. Consequently, the output of the discriminator can measure how closely synthetic data resembles real data. The performance of the discriminator on the synthesized data can be used as an indicator of how well the generator produces realistic data. This strategy can be applied not only to GANs but also to other generative models where a pre-trained machine learning model is used for evaluation.